40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16442     Entry time: Mon Nov 1 14:51:34 2021
Author: Koji 
Type: Update 
Category: General 
Subject: Checking the vent plan 

The vent team described a detailed vent plan (and reports where the actions have been performed)

https://wiki-40m.ligo.caltech.edu/vent/Fall2021

- [Sec.4] We should decide the final PR2 mirror through table-top measurements.

- [Sec 6] BS alignment is probably "unknown" now. So it'd be better to use the ITMY spot as the reference, then align BS for ITMX. For temporary alignment, it's OK though.

- [Sec 9-11] RIght now there is no mounts to place LO3/LO4/AS2/AS3/BHDBS. But we probably want to test something before the installation of the BHD? Just place the BHDBS on a optics mount so that we get an interfered beam on ITMY?

At this point we are supposed to have all the electronics all the CDS necessary for the new SOS control. Otherwise, they are just swinging and the alignment work will just be impossible.

- [Sec 15] The OPLEV mirrors can be freely moved as long as it does not block the main IR beams. Moving ITMXOL1 makes the reflection blocked by ITMXOL2. And moving ITMXOL2 would make the IR beams clipped. Consider replacing the mounts with a fixed mount. (The OPLEV mirrors are 1.5" in dia. It is not common vacuum compatible 1.5inch mounts. If 1" Al mirror is sufficient, we can use it.

https://wiki-40m.ligo.caltech.edu/vent/Fall2021/FinalAlignment

- The arms are the most strict alignment requirement. Everything else will follow the arm alignment. So start from the arms and propagate the alignment to Michelson / PRMI / SRMI.

- We reestablish arm alignment using the end green beams.

- Then recover IR arm alignment. Consider using ASS if possible

ELOG V3.1.3-